# Joyce structures from quadratic differentials on the sphere

### v1

@@ Acknowledgements
$\mathcal{P}(x)=\mathcal{F}_{I}(x)\Bigg{(}\frac{3}{4}\hbar^{2}+\hbar p_{I}(x-q_{I})+p_{I}^{2}(x-q_{I})^{2}+R_{3}(x)(x-q_{I})^{3}\Bigg{)}$ (B.4)
where $R_{3}(x)$ is a rational function in $x$ regular at $q_{I}$ with $R_{3}(q_{I})\neq 0$.
Suppose that $\mathcal{P}(x)$ had a repeated zero for all $\hbar$ with $|\hbar|<\epsilon$. Well then for small enough $\epsilon$ and for some $I$ there exists a holomorphic function $a_{I}:B(0,\epsilon)\to\mathbb{C}$ of $\hbar$ with $a_{I}(0)=q_{I}$ and such that $\mathcal{P}(a_{I}(\hbar))=0$ and $\mathcal{P}^{\prime}(a_{I}(\hbar))=0$. Since we can take $\epsilon$ small enough so $\mathcal{F}_{I}(a_{I}(\epsilon))\neq 0$ we obtain from $\mathcal{P}(a_{I}(\hbar))=0$ and $\mathcal{P}^{\prime}(a_{I}(\hbar))=0$ respectively:
$\frac{3}{4}\hbar^{2}+\hbar p_{I}(a_{I}(\hbar)-q_{I})+p_{I}^{2}(a_{I}(\hbar)-q_{I})^{2}+R_{3}(a_{I}(\hbar))(a_{I}(\hbar)-q_{I})^{3}$ $=0$ (B.5)
$\hbar p_{I}+2p_{I}^{2}(a_{I}(\hbar)-q_{I})+3R_{3}(a_{I}(\hbar))(a_{I}(\hbar)-q_{I})^{2}+R_{3}^{\prime}(a_{I}(\hbar))(a_{I}(\hbar)-q_{I})^{3}$ $=0.$ (B.6)
Eliminating the $(a_{I}(\hbar)-q_{I})^{3}$ and $(a_{I}(\hbar)-q_{I})^{0}$ terms respectively, yields the two equations
$\frac{9}{4}+2p_{I}\frac{(a_{I}(\hbar)-q_{I})}{\hbar}+p_{I}^{2}\frac{(a_{I}(\hbar)-q_{I})^{2}}{\hbar^{2}}-R^{\prime}_{3}(a_{I}(\hbar))\frac{(a_{I}(\hbar)-q_{I})^{4}}{\hbar^{2}}=0$ (B.7)
@@
$-p_{I}^{2}\frac{(a_{I}(\hbar)-q_{I})}{\hbar}-R_{3}(a_{I}(\hbar))\frac{(a_{I}(\hbar)-q_{I})}{\hbar}(a_{I}(\hbar)-q_{I})\Bigg{)}=0.$ (B.8)
Now
$\lim_{\hbar\to 0}\frac{a_{I}(\hbar)-q_{I}}{\hbar}=a_{I}^{\prime}(0)$ (B.9)
@@
$\frac{9}{4}+2p_{I}a_{I}^{\prime}(0)+p_{I}^{2}a_{I}^{\prime}(0)^{2}$ $=0$ (B.10)
$a_{I}^{\prime}(0)\bigg{(}\frac{1}{2}p_{I}-p_{I}^{2}a_{I}^{\prime}(0)\bigg{)}$ $=0.$ (B.11)
@@
Appendix C C. Linear independence of isomonodromic flows
@@
given in Appendix A are linearly independent, hence trivialise $TX$. This would then complete the proof of Proposition 7.3 by showing that the basis of isomonodromic flows $\{L_{i}^{(\alpha)}\}\cup\{L_{I}\}$ define a twistor distribution of rank $2n$ in the sense of Definition 2.2.
First, note that the linear independence of the $n$ vector fields $\{V_{i}^{(\alpha)}\}$ is equivalent to the invertibility of a particular $n\times n$ matrix $\mathcal{N}$ (related to a Vandermonde matrix) with entries $\mathcal{N}^{I}_{(\alpha,i)}$, letting $(\alpha,i)$ index the column and $I$ index the row
$\mathcal{N}_{(\alpha,i)}^{I}\coloneqq 2p_{I}\frac{\partial p_{I}}{\partial a_{i}^{(\alpha)}}=\frac{1}{(q_{I}-w_{\alpha})^{i}}$ (C.1)
@@
$\mathcal{N}_{(\infty,i)}^{I}\coloneqq p_{I}\frac{\partial p_{I}}{\partial a_{i}^{(\infty)}}=q_{I}^{i}$ (C.2)
for $i=0,...,m_{\infty}-4$, I=1,,nI=1,...,n. We will see later $\operatorname{det}(\mathcal{N)}\neq 0$ on $X$.
Given this non-vanishing we see that the $v_{I}$ coordinate derivatives for I=1,,nI=1,...,n are in the span of $\mathcal{S}$. We can thus instead consider a set $\tilde{S}$ with $\text{span}(\tilde{S})=\text{span}(S)$ given by
$\tilde{S}=\{\tilde{U_{I}},\tilde{V_{I}}\}\cup\{U_{i}^{(\alpha)},V_{i}^{(\alpha)}\}$ (C.3)
@@
$\Bigg{\{}\sum_{i=m_{\infty}-3}^{2m_{\infty}-7}\hskip-5.69054ptT_{i}^{(\infty)}(q_{I})\frac{\partial}{\partial a_{i}^{(\infty)}}+\sum_{\alpha=1}^{n}\sum_{i=m_{\alpha}+1}^{2m_{\alpha}-1}\hskip-5.69054ptT_{i}^{(\alpha)}(q_{I}-w_{\alpha})\frac{\partial}{\partial{a_{i}^{(\alpha)}}}+\sum_{\alpha=1}^{N}\frac{1}{q_{I}-w_{\alpha}}\frac{\partial}{\partial a_{2m_{\alpha}}^{(\alpha)}}\Bigg{\}}_{I=1}^{n}$ (C.4)
@@
$\Bigg{\{}\frac{\partial}{\partial{a_{i}^{(\alpha)}}}\Bigg{\}}_{\alpha=1,...,N}^{i=m_{\alpha}+1,...,2m_{\alpha}}\cup\Bigg{\{}\frac{\partial}{\partial a_{i}^{(\infty)}}\Bigg{\}}^{i=m_{\infty}-3,...,2m_{\infty}-7}$ (C.5)
is invertible if and only if $\mathcal{N}$ is. This follows because we may express $\hat{\mathcal{N}}=\mathcal{N}\mathcal{T}$, where $\mathcal{T}$ is block diagonal, with $N+1$, $m_{\alpha}\times m_{\alpha}$ blocks labelled by $\alpha=1,...,N,\infty$ corresponding to the poles. Each block is triangular with non-vanishing diagonal. The entries of $\mathcal{T}$ comes from the coefficients of the functions $T_{i}^{(\alpha)}$ and $T_{i}^{(\infty)}$ written down in Appendix A. Schematically, for $\alpha=1,...,N$ we have
$\begin{bmatrix}\frac{1}{q_{1}-w_{\alpha}}&\frac{1}{(q_{1}-w_{\alpha})^{2}}&\cdots&\frac{1}{(q_{1}-w_{\alpha})^{m_{\alpha}}}\\ \frac{1}{q_{2}-w_{\alpha}}&\frac{1}{(q_{2}-w_{\alpha})^{2}}&\cdots&\frac{1}{(q_{2}-w_{\alpha})^{m_{\alpha}}}\\ \vdots&\vdots&\vdots&\vdots\\ \frac{1}{q_{n}-w_{\alpha}}&\frac{1}{(q_{n}-w_{\alpha})^{2}}&\cdots&\frac{1}{(q_{n}-w_{\alpha})^{m_{\alpha}}}\end{bmatrix}\hskip-8.53581pt\begin{bmatrix}1&0&\cdots&0\\ 0&{(2m_{\alpha}-3)a_{2m_{\alpha}-1}^{(\alpha)}}&\cdots&(m_{\alpha}-1)a_{m_{\alpha}+1}\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&(2m_{\alpha}-3)(m_{\alpha}-1)!a_{2m_{\alpha}-1}\end{bmatrix}$
$=\begin{bmatrix}\frac{1}{q_{1}-w_{\alpha}}&T^{(\alpha)}_{2m_{\alpha}-1}(q_{1}-w_{\alpha})&\cdots&T^{(\alpha)}_{m_{\alpha}+1}(q_{1}-w_{\alpha})\\ \frac{1}{q_{2}-w_{\alpha}}&T^{(\alpha)}_{2m_{\alpha}-1}(q_{2}-w_{\alpha})&\cdots&T^{(\alpha)}_{m_{\alpha}+1}(q_{2}-w_{\alpha})\\ \vdots&\vdots&\vdots&\vdots\\ \frac{1}{q_{n}-w_{\alpha}}&T^{(\alpha)}_{2m_{\alpha}-1}(q_{n}-w_{\alpha})&\cdots&T^{(\alpha)}_{m_{\alpha}+1}(q_{n}-w_{\alpha})\\ \end{bmatrix}$, (C.6)
@@
So it remains to check that $\mathcal{N}$ is invertible on $X$. Recall that on $X$ the $q_{I}$ are distinct and away from the branch points. The poles $w_{\alpha}$ are of course also distinct. Given these conditions we will show $\det(\mathcal{N})$ is non-vanishing.
@@
$\mathcal{M}=\prod_{\begin{subarray}{c}\alpha=1,...,N\\ I=1,...,n\end{subarray}}(w_{\alpha}-q_{I})^{m_{\alpha}}\mathcal{N}.$ (C.7)
@@
$=n^{2}(n-m_{\infty}+3)-\frac{n}{2}+\frac{(m_{\infty}-3)^{2}}{2}-\sum_{\alpha=1}^{N}\frac{m_{\alpha}^{2}}{2},$ (C.8)
@@
$\frac{d}{dt}\det(\mathcal{N})=\pm\hskip-8.53581pt\sum_{i=1,..,m_{\alpha}}(-1)^{i}\det(\mathcal{N}{[\alpha,i]})$ (C.9)

### v2

@@ Acknowledgements
$\mathcal{P}(x)=\mathcal{F}_{I}(x)\Bigg(\frac{3}{4}\hbar^{2}+\hbar p_{I}(x-q_{I})+p_{I}^{2}(x-q_{I})^{2}+R_{3}(x)(x-q_{I})^{3}\Bigg),$ (C.4)
where $R_{3}(x)$ is a rational function in $x$ regular at $q_{I}$.
Suppose that $\mathcal{P}(x)$ had a repeated zero for all $\hbar$ with $|\hbar|<\epsilon$. Then, for small enough $\epsilon$ and for some $I$ there exists a holomorphic function $a_{I}:B(0,\epsilon)\to\mathbb{C}$ of $\hbar$ with $a_{I}(0)=q_{I}$ and such that $\mathcal{P}(a_{I}(\hbar))=0$ and $\mathcal{P}^{\prime}(a_{I}(\hbar))=0$. Since we can take $\epsilon$ small enough so $\mathcal{F}_{I}(a_{I}(\epsilon))\neq 0$, we obtain from $\mathcal{P}(a_{I}(\hbar))=0$ and $\mathcal{P}^{\prime}(a_{I}(\hbar))=0$ respectively:
$\frac{3}{4}\hbar^{2}+\hbar p_{I}(a_{I}(\hbar)-q_{I})+p_{I}^{2}(a_{I}(\hbar)-q_{I})^{2}+R_{3}(a_{I}(\hbar))(a_{I}(\hbar)-q_{I})^{3}$ $=0,$ (C.5)
$\hbar p_{I}+2p_{I}^{2}(a_{I}(\hbar)-q_{I})+3R_{3}(a_{I}(\hbar))(a_{I}(\hbar)-q_{I})^{2}+R_{3}^{\prime}(a_{I}(\hbar))(a_{I}(\hbar)-q_{I})^{3}$ $=0.$ (C.6)
Eliminating the $(a_{I}(\hbar)-q_{I})^{3}$ and $(a_{I}(\hbar)-q_{I})^{0}$ terms respectively yields the two equations
$\frac{9}{4}+2p_{I}\frac{(a_{I}(\hbar)-q_{I})}{\hbar}+p_{I}^{2}\frac{(a_{I}(\hbar)-q_{I})^{2}}{\hbar^{2}}-R^{\prime}_{3}(a_{I}(\hbar))\frac{(a_{I}(\hbar)-q_{I})^{4}}{\hbar^{2}}=0$ (C.7)
@@
$-p_{I}^{2}\frac{(a_{I}(\hbar)-q_{I})}{\hbar}-R_{3}(a_{I}(\hbar))\frac{(a_{I}(\hbar)-q_{I})}{\hbar}(a_{I}(\hbar)-q_{I})\Bigg)=0.$ (C.8)
Now,
$\lim_{\hbar\to 0}\frac{a_{I}(\hbar)-q_{I}}{\hbar}=a_{I}^{\prime}(0)$ (C.9)
@@
$\frac{9}{4}+2p_{I}a_{I}^{\prime}(0)+p_{I}^{2}a_{I}^{\prime}(0)^{2}$ $=0,$ (C.10)
$a_{I}^{\prime}(0)\bigg(\frac{1}{2}p_{I}-p_{I}^{2}a_{I}^{\prime}(0)\bigg)$ $=0.$ (C.11)
@@
Appendix D D. Linear independence of isomonodromic flows
@@
given in Appendix B are linearly independent, hence trivialise $TX$. This would then complete the proof of Proposition 7.3 by showing that the basis of isomonodromic flows $\{L_{i}^{(\alpha)}\}\cup\{L_{I}\}$ define a twistor distribution of rank $2n$ in the sense of Definition 2.2.
First, note that the linear independence of the $n$ vector fields $\{V_{i}^{(\alpha)}\}$ is equivalent to the invertibility of a particular $n\times n$ matrix $\mathcal{N}$ (related to a Vandermonde matrix) with entries $\mathcal{N}^{I}_{(\alpha,i)}$, letting $(\alpha,i)$ index the column and $I$ index the row:
$\mathcal{N}_{(\alpha,i)}^{I}\coloneqq 2p_{I}\frac{\partial p_{I}}{\partial a_{i}^{(\alpha)}}=\frac{1}{(q_{I}-w_{\alpha})^{i}}$ (D.1)
@@
$\mathcal{N}_{(\infty,i)}^{I}\coloneqq 2p_{I}\frac{\partial p_{I}}{\partial a_{i}^{(\infty)}}=q_{I}^{i}$ (D.2)
for $i=0,...,m_{\infty}-4$, I=1,,nI=1,...,n.
We will see later $\operatorname{det}(\mathcal{N)}\neq 0$ on $X$. Given this non-vanishing, we see that the $v_{I}$ coordinate derivatives for I=1,,nI=1,...,n are in the span of $\mathcal{S}$. We can thus instead consider a set $\tilde{\mathcal{S}}$ with $\text{span}(\tilde{\mathcal{S}})=\text{span}(\mathcal{S})$ given by
$\tilde{S}=\{\tilde{U_{I}},\tilde{V_{I}}\}\cup\{U_{i}^{(\alpha)},V_{i}^{(\alpha)}\},$ (D.3)
@@
$\Bigg\{\sum_{i=m_{\infty}-3}^{2m_{\infty}-7}\hskip-5.69054ptT_{i}^{(\infty)}(q_{I})\frac{\partial}{\partial a_{i}^{(\infty)}}+\sum_{\alpha=1}^{n}\sum_{i=m_{\alpha}+1}^{2m_{\alpha}-1}\hskip-5.69054ptT_{i}^{(\alpha)}(q_{I}-w_{\alpha})\frac{\partial}{\partial{a_{i}^{(\alpha)}}}+\sum_{\alpha=1}^{N}\frac{1}{q_{I}-w_{\alpha}}\frac{\partial}{\partial a_{2m_{\alpha}}^{(\alpha)}}\Bigg\}_{I=1}^{n}$ (D.4)
@@
$\Bigg\{\frac{\partial}{\partial{a_{i}^{(\alpha)}}}\Bigg\}_{\alpha=1,...,N}^{i=m_{\alpha}+1,...,2m_{\alpha}}\cup\Bigg\{\frac{\partial}{\partial a_{i}^{(\infty)}}\Bigg\}^{i=m_{\infty}-3,...,2m_{\infty}-7}$ (D.5)
is invertible if and only if $\mathcal{N}$ is. This follows because we may express $\hat{\mathcal{N}}=\mathcal{N}\mathcal{T}$, where $\mathcal{T}$ is block diagonal, with $N+1$, $m_{\alpha}\times m_{\alpha}$ blocks labelled by $\alpha=1,...,N,\infty$ corresponding to the poles. Each block is triangular with non-vanishing diagonal. The entries of $\mathcal{T}$ comes from the coefficients of the functions $T_{i}^{(\alpha)}$ and $T_{i}^{(\infty)}$ written down in Appendix B. Schematically, for $\alpha=1,...,N$ we have
$\begin{bmatrix}\frac{1}{q_{1}-w_{\alpha}}&\frac{1}{(q_{1}-w_{\alpha})^{2}}&\cdots&\frac{1}{(q_{1}-w_{\alpha})^{m_{\alpha}}}\\ \frac{1}{q_{2}-w_{\alpha}}&\frac{1}{(q_{2}-w_{\alpha})^{2}}&\cdots&\frac{1}{(q_{2}-w_{\alpha})^{m_{\alpha}}}\\ \vdots&\vdots&\vdots&\vdots\\ \frac{1}{q_{n}-w_{\alpha}}&\frac{1}{(q_{n}-w_{\alpha})^{2}}&\cdots&\frac{1}{(q_{n}-w_{\alpha})^{m_{\alpha}}}\end{bmatrix}\hskip-8.53581pt\begin{bmatrix}1&0&\cdots&0\\ 0&{(2m_{\alpha}-3)a_{2m_{\alpha}-1}^{(\alpha)}}&\cdots&(m_{\alpha}-1)a_{m_{\alpha}+1}\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&{a_{2m_{\alpha}-1}}\end{bmatrix}$
$=\begin{bmatrix}\frac{1}{q_{1}-w_{\alpha}}&T^{(\alpha)}_{2m_{\alpha}-1}(q_{1}-w_{\alpha})&\cdots&T^{(\alpha)}_{m_{\alpha}+1}(q_{1}-w_{\alpha})\\ \frac{1}{q_{2}-w_{\alpha}}&T^{(\alpha)}_{2m_{\alpha}-1}(q_{2}-w_{\alpha})&\cdots&T^{(\alpha)}_{m_{\alpha}+1}(q_{2}-w_{\alpha})\\ \vdots&\vdots&\vdots&\vdots\\ \frac{1}{q_{n}-w_{\alpha}}&T^{(\alpha)}_{2m_{\alpha}-1}(q_{n}-w_{\alpha})&\cdots&T^{(\alpha)}_{m_{\alpha}+1}(q_{n}-w_{\alpha})\\ \end{bmatrix}$, (D.6)
@@
So it remains to check that $\mathcal{N}$ is invertible on $X$. Recall that on $X$ the $q_{I}$ are distinct and away from the branch points. The poles $w_{\alpha}$ are of course also distinct. Given these conditions, we will show $\det(\mathcal{N})$ is non-vanishing.
@@
$\mathcal{M}=\prod_{\begin{subarray}{c}\alpha=1,...,N\\ I=1,...,n\end{subarray}}(w_{\alpha}-q_{I})^{m_{\alpha}}\mathcal{N}.$ (D.7)
@@
$=n^{2}(n-m_{\infty}+3)-\frac{n}{2}+\frac{(m_{\infty}-3)^{2}}{2}-\sum_{\alpha=1}^{N}\frac{m_{\alpha}^{2}}{2},$ (D.8)
@@
$\frac{d}{dt}\det(\mathcal{N})=\pm\hskip-8.53581pt\sum_{i=1,..,m_{\alpha}}(-1)^{i}\det(\mathcal{N}{[\alpha,i]}),$ (D.9)
@@
Data availability
No datasets were generated or analysed during this work.
Conflicts of interest
The author declares there are no conflicts of interest.